OpenStudy (anonymous):

|x + 3| = 10 solve for x

OpenStudy (anonymous):

Do you know how to solve for absolute value?

OpenStudy (anonymous):

What Is 10 - 3?

OpenStudy (anonymous):

It is asking you, what plus 3 will give me 10

OpenStudy (anonymous):

This equation is absolute value. |x+3|=10 Step 1:Take away the absolute value lines, x+3=10 Step 2: Then solve the equation like a normal problem x+3=10 -3 -3 x=7 This is one of your solutions, the other solution is the opposite of what I just did. |x+3|=10 +3 +3 x=-13 (because this is opposite) Therefore your answers are x= 7,-13

OpenStudy (anonymous):

put like terms together so x= 10-3 which= 7 hence x=7